OpenStudy (anonymous):

write cot 0 in terms of cos 0

OpenStudy (homeworksucks):

I'm going to assume that you mean \[\theta\]

OpenStudy (anonymous):

\[ \cot(\theta) = \frac{\cos(\theta)}{\sin(\theta)} \]

OpenStudy (anonymous):

\[ \sin^2(\theta)+\cos^2(\theta) = 1 \implies |\sin(\theta)| = \sqrt{1-\cos^2(\theta)} \]

OpenStudy (homeworksucks):